OpenStudy (anonymous):

write an equation in point-slope form of the line that passes through the given points (6,-2), (12,1)?

OpenStudy (mertsj):

OK. After i showed you the last one, did you even try this one?

OpenStudy (mertsj):

It is very hard to help you if you won't answer me.

OpenStudy (mertsj):

Find the slope. Can you do that?

OpenStudy (mertsj):

Ok. Luis wants to help you. I'll turn you over to him.

OpenStudy (mertsj):

I see Luis has lost interest. Can you find the slope?

OpenStudy (anonymous):

hold my computers is messed up

OpenStudy (msmr):

First you find the slope: y2-y1/x2-x1

OpenStudy (msmr):

in this case, it would be 1-(-2)/12-6 or 3/6 -> 1/2

OpenStudy (msmr):

and then you pick a point and insert it into the point slope formula: y-12 = 1/2(x-1)

OpenStudy (anonymous):

is this the answer y-6=12(x+2)